OpenStudy (anonymous):

simplify 2 root a - 3 root 4a

OpenStudy (wolf1728):

Is this what you mean? 

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

Okay just making sure.

OpenStudy (wolf1728):

-3 *sqrt(4a) = -3*2* sqrt(a) = -6 * sqrt(a)

OpenStudy (wolf1728):

2 * sqrt(a) - 6* sqrt(a) = 4 * sqrt(a)

OpenStudy (anonymous):

thanks :)

OpenStudy (wolf1728):

u r welcome